# ReMatch: Retrieval Enhanced Schema Matching with LLMs

### v1

@@ 4.1. Dataset Creation
Table 1. Dataset statistics for MIMIC and Synthea. #Columns and #Tables refer to the number of columns and tables used in the dataset. #Mapped Columns refers to the total number of mappings for the source schemas, and in the case of target schemas (OMOP), the number of unique columns mapped. #Null Mappings refers to columns without a mapping (mapped to NA).
@@
4.2. Experiments
We established the performance of ReMatch and previous state-of-the-art (SOTA) methods on MIMIC and Synthea. In this process we aimed to both assess the abilities of ReMatch, as well as compare it to methods that use labeled data.
It’s important to note that while ReMatch was evaluated on the full mapping, including nulls, the evaluation of the other models focused solely on attributes with existing matches. This means that input attributes with no match in the target were not scored, which is a significant relaxation of the problem.
Prior methods, as explained in Section 5, approached schema matching as a binary classification problem, with the usual F1-score. Since our method treats it as a retrieval problem, the most appropriate metric is accuracy@$K$, as defined in Section 3.1. Unlike the F1-score, standard accuracy@$K$ is not well defined for m:n matches. To deal with this, we limited ourselves to evaluating 1:1 or m:1 relations. Nonetheless, it is worth noting that for a dataset with only 1:1 and m:1 mappings, accuracy@1 is equivalent to F1-score when using argmax, instead of a threshold, for prediction (since it trivially reflects both the precision and the recall simultaneously).
It is also important to note that there are multiple dimensions we would like to optimize simultaneously. The first is maximizing accuracy@$K$, and is the primary goal. We also seek to achieve better results with lower values of $K$, as our aim is to assist human schema matching, and minimize cognitive overload (Bernstein et al., 2011; Chernev et al., 2015). Additionally, we would also like to minimize the number of target tables, to reduce prompt size and therefore costs. Although choosing $J$ does not guarantee the number of target tables in the final prompt, minimizing $J$ does serve as a proxy for smaller prompts.
Evaluating ReMatch  For all experiments we used GPT-4 (OpenAI, 2023) as our generative LLM and Ada2 (Greene et al., 2022; Neelakantan et al., 2022) as our embedding model. The full prompts and hyperparameters can be found in Appendix A.
We performed a grid search over various values of $J$ and $K$, to establish the performance of our method. This enabled us to select hyperparameters for further evaluations.
We performed an ablation study regarding the usefulness of the first and second stage in ReMatch. We evaluated the performance on MIMIC when no descriptions are used, meaning that instead of using stage (1) in ReMatch, only the names of the tables and attributes are used for retrieval and generation. We also evaluated a version with no retrieval stage, i.e., for each table in the source schema we used the entire target schema.
Finally, we ran ReMatch multiple times on each of the chosen setups in the ablation study. This was done to deal with the stochastic nature of LLM text generation and to ensure the results are consistent.
Evaluating ReMatch with Guidance  In real-world mapping scenarios a human matcher may receive ReMatch’s output and determine that the suggestions are not good enough. Alternatively, a human matcher may decide to provide information they already have about matches. In either of these cases, ReMatch should allow the matcher to influence the output of the mechanism. This includes highlighting the relevancy of a certain table, or explicitly providing the mapping of an attribute.
In our evaluation we investigated a mechanism, inspired by the relevance feedback technique (Rocchio, 1971), of the impact this may have on ReMatch’s performance. To do so, we selected a mapping of the form $(T_{1},a_{1})\rightarrow(T_{2},a_{2})$ for a single attribute from each table. Specifically, we used the SUBJECT_ID attribute where available (19/25 tables), and an attribute containing unique identifiers otherwise. We then automatically included $T_{2}$ in the set of retrieved tables, $\mathcal{T}_{c}$, and provided the LLM with this mapping in the prompt.
Evaluating Other Models  Our method’s performance was evaluated against Ditto (Li et al., 2020) and SMAT (Zhang et al., 2021), two previously SOTA non-LLM models. For Ditto training we used the default hyperparameters: {Learning Rate: 3e-5, Batch Size: 64, Epochs: 20}, with early stopping. For SMAT training we also used the default hyperparameters: {Learning Rate: 0.8, Batch Size: 64, Epochs: 30}, with early stopping. In all setups, 10% of the training data was used as a validation set to avoid over-fitting. Models were trained on a single Nvidia A-100 GPU.
Both models demand a preprocessing step to generate a Cartesian product of all source and target attributes, and each pair is then labeled as positive for a match and negative otherwise. To compare ReMatch to these models we created a similar representation for our dataset (MIMIC) with a Cartesian product representation, which resulted in a positive to negative class ratio of approximately 1:400. While we utilized both MIMIC and Synthea to directly compare to ReMatch, we also used Synthea as a sanity check for verifying our ability to reproduce models at least as good as Ditto and SMAT’s reported results.
We used different split ratios of training and test subsets to understand how the models perform under varying conditions and prerequisites. Namely, we tried two setups, 80% training and 20% testing, and vice versa. The splits were stratified across table names. This approach also allows for an understanding of how much preliminary work is necessary to achieve similar scores to ReMatch on the MIMIC dataset.
We also extended the evaluation to include accuracy@$K$. To do so we adapted the models’ output to calculate rankings rather than binary predictions. We calculated the accuracy@$K$ by determining whether the correct match received one of the $K$ highest classification scores. Binary predictions for F1-score calculation were given by using the optimal threshold for the validation set.
@@ 4.3. Results
ReMatch Results The results of the grid search for MIMIC can be found in Table 2. As explained in Section 4.2, the highest value in a given column or row is not necessarily the optimal choice. Rather, we found $\{(J=1,\>K=1),\>(J=2,\>K=5)\}$ to be the best balance points between the three optimization dimensions for the MIMIC dataset. Note the strong, albeit non-linear, correlation between $J$ and the average number of tables (Avg #T) retrieved in practice. This is in contrast to accuracy@$K$, which seems to perform best for $J\leq 2$. This may imply that the retrieval is efficient, capturing all relevant tables at low values of $J$, and additional tables only add ‘noise’ to the matching prompt.
The results for the grid search for Synthea are presented inTable 3. While $\{(J=1,\>K=1),\>(J=2,\>K=5)\}$ still appear to work well, for Synthea we found that not using retrieval yielded optimal results for accuracy@5. This may be explained by the small size of the dataset, with only 8 tables in the target. It is still interesting to note that ReMatch achieved almost 75% accuracy@5 also with $J=2$. While the results are a bit higher for Synthea, the gap in performance between datasets is relatively small.
The results of the ablation study are shown in Table 4. We found that skipping the retrieval step and adding all of the tables yields inferior results, in alignment with the results from the grid search. We also found that using only the names of the tables and attributes, with no descriptions, results in a $6.5\%$ decrease in accuracy@1 and a $31\%$ decrease in accuracy@5. A possible explanation for this may be that ‘easy matches’ can be done based on name semantics alone, but more complex, less likely, matches require additional information. Additionally, Table 4 demonstrates that although ReMatch uses generative modeling, the method’s outputs are consistent.
@@
SMAT and Ditto Results  SMAT’s performance was successfully validated using the F1-score, yielding an F1-score of 0.48276 for the 80-20 train-test split on the Synthea dataset, indicating successful model training when compared to the original values reported by Zhang et al. The SMAT model outperformed the Ditto model across all metrics and setups. Findings from the Ditto model revealed poor performance in all schema matching setups. Notably, in all scenarios, accuracy@1 was $\leq 0.05$.
@@
Table 6. Results for Ditto and SMAT on both datasets.
@@
ptACC@1ptACC@5000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.8000.0610.0610.0610.0610.0770.0770.0770.0770.3610.3610.3610.3610.4240.4240.4240.4240.7290.7290.7290.729DittoSMATReMatch
(c) Performance on MIMIC.
ptACC@1ptACC@5000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811110.020.020.020.020.120.120.120.120.2710.2710.2710.2710.6940.6940.6940.6940.5620.5620.5620.5620.9240.9240.9240.924DittoSMATReMatch
(d) Performance on Synthea.
Figure 2. Comparison of the different models’ performance. Ditto and SMAT were trained and evaluated on 20%, 80% of the data, respectively, after removing all null mappings. ReMatch was evaluated on the entire dataset, with no guidance, and with nulls. Optimal setup from grid search is shown for ReMatch.
5. Related Work
Prior work has dealt with various ML based matchers to tackle the schema matching task, but these have often resulted in low-quality matches or necessitated extensive manual annotation (Bernstein et al., 2011).
More recent mechanisms often employed similarity matrices between schemas, with various ML or deep learning (DL) components (Mudgal et al., 2018; Shraga et al., 2020; Li et al., 2020), but these were limited to small datasets, like Purchase Orders (Do and Rahm, 2002), or Web Forms (Gal, 2011), and focused primarily on entity matching rather than schema matching. More recent work in the field such as SMAT (Zhang et al., 2021) and LSM (Zhang et al., 2023b) has been proposed, leveraging recent improvements in natural language processing techniques to obtain semantic mappings between source and target schemas. These approaches improved previous results, but still demanded that a large percentage of the data be tagged, making real-world usage impractical.
LLMs have been applied to various data preprocessing tasks, including error detection, data imputation, entity matching, as well as schema matching (Mirchandani et al., 2023; Narayan et al., 2022). These models have shown promising results, with LLMs achieving good results on various datasets, with no additional training involved. However, these approaches still faced challenges related to computational expense and inefficiency, as well as lacking evaluation on data that represents real-world schema matching challenges.
Finally, it is important to note that the majority of previous methods dealt with schema-matching as a binary classification task over the full Cartesian-product $\mathcal{A}_{1}\times\mathcal{A}_{2}$. While allowing for simplified training objectives and evaluations, in normalized schemas positive labels scale by $O(n)$, whereas negative labels scale by $O(n^{2})$. This is especially problematic when inference is expensive, as in the case of LLMs.
6. Discussion
Schema matching is a complex task challenged by heterogeneity across schemas, semantic ambiguity of schema elements, and the scale and complexity of large enterprise systems. Accurate matching is further complicated by the need for domain-specific knowledge, limited context for interpreting schema elements, variability in database design practices, and data privacy concerns that may restrict access to the actual data.
Our evaluation has shown that ReMatch achieves good accuracy results in practice. This was done with no labels being provided, on a large schema matching dataset, while maintaining efficiency in prompt size. We found that ReMatch can be further improved by incorporating additional information in the form of guidance. Additionally, we have shown that other methods indeed demand significant amounts of annotated data to achieve good performance, demonstrating the impracticality of direct usage in real-world scenarios.
Limitations  While the ReMatch method presents an innovative approach to schema matching, and has demonstrated significant improvements in matching capabilities and outperforms other machine learning approaches, it is important to acknowledge its limitations.
Similar to human matchers, the effectiveness of ReMatch is contingent upon the quality of schema documentation. The approach relies on the presence of meaningful labels and textual descriptions for tables and attributes within the schema. In cases where schema documentation is poor or lacks clarity, ReMatch might struggle to perform optimally.
While we have successfully demonstrated high accuracy using the ReMatch method, we acknowledge that we have not yet delved into the influence of varying document structures and prompt designs on the matching results. The current implementation utilizes a specific document format and prompt configuration, which has proven to be effective; however, it is conceivable that alternative configurations may further optimize performance. This aspect represents an area for potential refinement rather than a significant drawback. Future investigations into these variables may provide incremental improvements and contribute to the fine-tuning of our approach, ensuring that ReMatch remains at the forefront of schema matching solutions.
@@ 7. Conclusions
In this work we introduced ReMatch, a scalable and efficient method for matching schemas using retrieval-enhanced LLMs. It starts by representing source schema attributes and target schema tables as structured documents. For each source attribute, the method retrieves the most semantically similar target table documents, constructing a candidate set for potential mapping. Subsequently, a prompt comprising these candidates is formulated, and an LLM is employed to identify the top matches for each source attribute.
ReMatch avoids the need for predefined mapping, any model training, or access to data in the source database. Instead, it exploits the generative abilities of LLMs, and their text comprehension to perform semantic ranking between two schemas, in alignment with a human matching process.
ReMatch is not designed to completely replace human matchers. It aims to complement their efforts by acting as a schema matching Copilot, minimizing human error and aiding matchers throughout their work.
Future Work  Adjusting the structure of documents and the prompts used with LLMs to better suit specific domains could significantly improve the accuracy and relevance of the matches produced by ReMatch. This could involve tailoring the language and context within the prompts to align with the specific jargon and data relationships found in each industry. By fine-tuning these aspects, the method can become more sensitive to the nuances of each industry’s data representation needs.
Moreover, in situations where the source schema data is available and can be accessed without privacy or security constraints, this data can be helpful for enrichment of labels and the potential for generating more informative labels in cases where existing documentation is sparse or ambiguous.
Finally, further research is needed regarding combining ReMatch with additional algorithmic improvements. These may include explicit usage of type constraints, foreign keys, and primary keys, both directly and as post-processing steps, as well as additional usage of guidance mechanisms. New methods combining ReMatch, for initial labeling of the majority of the data, with other methods for downstream inference, may be of interest. Such an approach could leverage the advantages of each method to achieve even better results.
Appendix A Full Matching Prompt
A.1. Prompt
You are an expert in databases, and schema matching at top k specifically. Your task is to create matches between source and target tables and attributes. For each attribute from the source you always suggest the top 2 most relevant tables and columns from the target. You are excellent at this task.
If none of the columns are relevant, the last table and column should be ”NA”, ”NA”. This value may appear only once per mapping!
Your job is to match the schemas. You never provide explanations, code or anything else, only results. Below are the two schemas.
Create top k matches between source and target tables and columns.
Make sure to match the entire input. Make sure to return the results in the following json format with top 2 target results foreach input in source.
Expected output format:
{’1’: {’SRC_ENT’: ’SOURCE_TABLE_NAME’, ’SRC_ATT’: ’SOURCE_COLUMN_NAME’,
’TGT_ENT1’: ’TARGET_TABLE_NAME1’, ’TGT_ATT1’: ’TARGET_COLUMN_NAME1’, ’TGT_ENT2’: ’TARGET_TABLE_NAME2’, ’TGT_ATT2’: ’TARGET_COLUMN_NAME2’},
’2’: {’SRC_ENT’: ’SOURCE_TABLE_NAME’, ’SRC_ATT’: ’SOURCE_COLUMN_NAME’,
’TGT_ENT1’: ’TARGET_TABLE_NAME1’, ’TGT_ATT1’: ’TARGET_COLUMN_NAME1’, ’TGT_ENT2’: ’TARGET_TABLE_NAME2’, ’TGT_ATT2’: ’TARGET_COLUMN_NAME2’}}
Source Schema:
,SRC_ENT, SRC_ATT
0,ADMISSIONS, SUBJECT_ID
1,ADMISSIONS, HADM_ID
[…]
ADMISSIONS
The ADMISSIONS table gives information regarding a patient’s admission to the hospital.
Primary Keys:
HADM_ID (INTEGER): Each row of this table contains a unique HADM_ID, which represents a single patient’s admission to the hospital. HADM_ID ranges from 1000000 - 1999999.
Foreign Keys:
SUBJECT_ID (INTEGER): The ADMISSIONS table can be linked to the PATIENTS table using SUBJECT_ID. References to [PATIENTS, SUBJECT_ID]
Other Columns:
ADMITTIME (TIMESTAMP): provides the date and time the patient was admitted to the hospital
[…]
Target Schema:
,TGT_ENT,TGT_ATT
0,PERSON,person_id
1,PERSON,gender_concept_id
[…]
19,VISIT_OCCURRENCE,visit_occurrence_id
20,VISIT_OCCURRENCE,person_id
[…]
PERSON
This table serves as the central identity management for all Persons in the database. It contains records that uniquely identify each person or patient, and some demographic information.
Primary Keys:
person_id (bigint): It is assumed that every person with a different unique identifier is in fact a different person and should be treated independently.
Foreign Keys:
gender_concept_id (integer): This field is meant to capture the biological sex at birth of the Person. This field should not be used to study gender identity issues. References to [CONCEPT, concept_id]
[…]
Remember to match the entire input. Make sure to return only the results!
A.2. GPT-4 Hyperparameters
The model version used for generation was GPT-4-1106, with the following settings:
{seed=42 temperature=0.5, max_tokens=4096, top_p=0.9, frequency_penalty=0, presence_penalty=0}

### v2

@@ 4.2. Experiments
Table 1. Dataset statistics for MIMIC and Synthea. #Columns and #Tables refer to the number of columns and tables used in the dataset. #Mapped Columns refers to the number of unique columns mapped. #Null Mappings refers to source columns without a mapping (mapped to NA).
@@ 4.3. Results
ReMatch Results The results of the grid search for MIMIC can be found in Table 2. We found $\{(J=1,\>K=1),\>(J=2,\>K=5)\}$ to be the best balance points between the three optimization dimensions for the MIMIC dataset. Interestingly, accuracy@$K$ seems to perform best for $J\leq 2$. This may imply that the retrieval is efficient, and additional tables only add ‘noise’ to the matching prompt.
The results for the grid search for Synthea are presented inTable 3. While $\{(J=1,\>K=1),\>(J=2,\>K=5)\}$ still appear to work well, for Synthea we found that not using retrieval yielded optimal results for accuracy@5. This may be explained by the small size of the dataset, with only 8 tables in the target.
The results of the ablation study are shown in Table 4. We found that skipping the retrieval step yields inferior results, in alignment with the results from the grid search. We also found that using only the names of the tables and attributes, with no descriptions, results in a significant decrease. Additionally, Table 4 demonstrates that the method’s outputs are stable.
@@
SMAT Results  SMAT’s performance was validated using the F1-score, with an F1-score of 0.48276 for the 80-20 train-test split on the Synthea dataset, indicating successful model training when compared to the original values reported by Zhang et al.
@@
ACC@1ACC@5000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.80.0770.0770.0770.0770.3610.3610.3610.3610.4240.4240.4240.4240.7290.7290.7290.729SMATReMatch
(a) Performance on MIMIC.
ACC@1ACC@5000.50.50.50.511110.2710.2710.2710.2710.6940.6940.6940.6940.5620.5620.5620.5620.9240.9240.9240.924SMATReMatch
(b) Performance on Synthea.
Figure 2. Comparison of the different models’ performance. SMAT was trained and evaluated on 20%, 80% of the data, after removing all null mappings. ReMatch was evaluated on the entire dataset, with no guidance, and with nulls. Optimal setup from grid search is shown for ReMatch.
Table 6. Results for SMAT on both datasets.
@@ 5. Conclusions
In this work we introduced ReMatch, a scalable and effective method for matching schemas using retrieval-enhanced LLMs, designed to complement and aid human matchers throughout their work. ReMatch avoids the need for predefined mapping, any model training, or access to data in the source database. Instead, it exploits the generative abilities of LLMs to perform semantic ranking between two schemas, in alignment with a human matching process. We also provided a new large dataset that will hopefully aid further research.
In future work, we plan to include the explicit use of type constraints, foreign keys, and primary keys, as well as enhanced guidance mechanisms and enrichment of table and column descriptions using the source schema data, where accessible without privacy or security constraints.